# DPA-Net: Structured 3D Abstraction from Sparse Views via Differentiable Primitive Assembly

### v2

@@ 3.3 Differentiable rendering
3.4 Network training and test-time adaption
Similar to [53, 25, 39], DPA-Net can work in a feed-forward fashion by training on a dataset including sparse-view images. We also propose a multi-phase test-time adaption for instance-specific refinement during inference.
@@ 3.4 Network training and test-time adaption
Test-time adaption (TTA): We have found that DPA-Net trained on a generic dataset risks suboptimal performance when applied to novel shapes which deviate substantially from the training data. For example, assembling each 3D object requires unique matrices $\mathbf{T}$ and $\mathbf{w}$ tailored to that instance. To improve generalization, we propose a test-time adaptation that operates in three phases with different configurations as shown in Table 1. Each phase is initialized from its preceding phase.
@@ 4 Network Improvements
Primitive dropouts: Attaining a compact primitive assembly, with a low primitive count, without compromising abstraction accuracy is clearly desirable. Similar to [54], we drop out a primitive if excluding it from the construction does not significantly alter the overall shape. In practice, we initialize with the primitives obtained from Phase 3. For each $p$th primitive, we set the $p$th row of the binary selection matrix $\mathbf{T}$ to zero, compute a new occupancy field $\mathbf{b}^{*}$ using Eq.1, and finally quantify the shape variation as: $v=\frac{1}{K}\sum_{i=1}^{K}\mathbbm{1}\left(\text{XOR}(\mathbf{b}_{i}^{*}<0.1,\mathbf{a}_{i}^{*}<0.1)\right),$ where $K=40,960$ is the number of points sampled around the shape as in Phase 3 and $\mathbbm{1}()$ maps Boolean values to 0 or 1.
@@ 5 Results & Evaluation
Metrics: We evaluate our method from two aspects. (1) 3D abstraction quality, measured by three standard metrics [6, 55]: symmetric Chamfer Distance (CD), Edge Chamfer Distance (ECD), and Normal Consistency (NC). The same parameter settings from CAPRI-Net [55] are used to calculate CD and ECD. (2) Shape compactness, which is reflected by having fewer parts (#Parts) and leads to ease of manipulation/editing.
@@ 5.1 Evaluation on ShapeNet
In particular, DPA-Net is better by 88%/17%/68% in CD/NC/ECD and uses 45% fewer 3D parts. Compared with the modified pixelNeRF, we obtain similar errors (differs by 0.3 in CD, 0.07 in NC, 0.8 in ECD) across all metrics. However, our method is more flexible in that it also produces convex part decomposition in an unsupervised manner, which is more useful in many applications, compared to the non-structured mesh produced by pixelNeRF.
@@
Method Input CD \!$\downarrow$ NC \!$\uparrow$ ECD \!$\downarrow$ #Parts \!$\downarrow$
EMS 3D 6.93 0.65 12.56 14.92
pixelNeRF 2D 0.50 0.85 3.21 -
Ours 0.79 0.78 4.01 8.15
Ablation studies: We validate the effectiveness of three important strategies on improving DPA-Net: adaptive pixel sampling (AS), overlapping loss (OL) and primitive dropout (PD). Table 3 shows the main quantitative results. Removing adaptive sampling (w/o AS) significantly degrades the abstraction quality as the CD increases by 29%, showing the importance of our sillouate-aware pixel sampling method. Removing the overlapping loss (w/o OL) results in higher number of primitives and convex shapes, showing that this loss helps in removing redundant, heavily overlapped convexes.
The primitive dropout (PD) strategy further reduces the number of primitives by 29% without sacrificing the abstraction accuracy. Fig. 4 qualitatively demonstrates the contributions of the three components. Adaptive pixel sampling helps recovering thin structures of the chair. Enforcing a overlap loss removes redundant 3D parts. Primitive dropouts effectively reduce the number of primitives for assembling the final shape.
Figure 4: Ablation studies on the three key strategies to boost DPA-Net. The texts, #P’s, in the last column show the number of qudaric primitives used. The cyan ovals highlight recovered thin structures or areas with holes. The red ovals highlight redundant overlapping convex parts.
Table 3: Ablation study on the ShapeNet chair benchmark. #P and #C denote the number of quadric primitives and convex shapes used during shape assembly. The colors cyan and orange represent the best and the second best results.
@@
More importantly, while pixelNeRF attained higher accuracy when trained solely on chairs, our DPA-Net surpasses pixelNeRF across all metrics in this diverse multi-category experiment — for instance, improving on the CD metric by nearly 60%. This highlights the enhanced capacity of our structured primitive-based representation in capturing expansive shape variations compared to pixelNeRF’s unstructured approach.
@@
Method Input CD \!$\downarrow$ NC \!$\uparrow$ ECD \!$\downarrow$ #Parts \!$\downarrow$
EMS 3D 5.15 0.75 43.45 8.28
pixelNeRF 2D 3.67 0.73 11.01 -
Ours 2D 1.47 0.70 4.91 5.57
@@ 5.2 Evaluation on DTU
As shown in Table 5, DPA-Net obtains higher-quality 3D abstractions with 24% lower average CD while using 3.5 more parts compared to DBW. DiViNet [46] and pixelNerf [53] outperform DAP-Net as they output un-structured meshes. Fig. 6 provides visual comparisons showing that our method can recover fine structural details.
@@
Method Input Structure Mean CD Mean #Parts
EMS 3D 4.37 4.50
DBW 2D 2.67 3.83
Ours 2D 2.04 7.33
pixelNerf 2D $\times$ 1.95 -
DiViNet 2D $\times$ 1.82 -

### v3

@@ 3.3 Differentiable rendering
3.4 Network training and test-time adaptation
Similar to [53, 25, 39], DPA-Net can work in a feed-forward fashion by training on a dataset including sparse-view images. We also propose a multi-phase test-time adaptation for instance-specific refinement during inference.
@@ 3.4 Network training and test-time adaptation
Test-time adaptation (TTA): We have found that DPA-Net trained on a generic dataset risks suboptimal performance when applied to novel shapes which deviate substantially from the training data. For example, assembling each 3D object requires unique matrices $\mathbf{T}$ and $\mathbf{w}$ tailored to that instance. To improve generalization, we propose a test-time adaptation that operates in three phases with different configurations as shown in Table 1. Each phase is initialized from its preceding phase.
@@ 4 Network Improvements
Primitive dropouts: Attaining a compact primitive assembly, with a low primitive count, without compromising abstraction accuracy is clearly desirable.
Similar to [54], we drop out a primitive if excluding it from the construction does not significantly alter the overall shape. In practice, we initialize with the primitives obtained from Phase 3. For each $p$th primitive, we set the $p$th row of the binary selection matrix $\mathbf{T}$ to zero, compute a new occupancy field $\mathbf{b}^{*}$ using Eq.1, and finally quantify the shape variation as: $v=\frac{1}{K}\sum_{i=1}^{K}\mathbbm{1}\left(\text{XOR}(\mathbf{b}_{i}^{*}<0.1,\mathbf{a}_{i}^{*}<0.1)\right),$ where $K=40,960$ is the number of points sampled around the shape as in Phase 3 and $\mathbbm{1}()$ maps Boolean values to 0 or 1.
@@ 5 Results & Evaluation
Metrics: We evaluate our method from three aspects. (1) 3D abstraction quality, measured by three standard metrics [6, 55]: symmetric Chamfer Distance (CD), Edge Chamfer Distance (ECD), and Normal Consistency (NC). The same parameter settings from CAPRI-Net [55] are used to calculate CD and ECD. (2) Shape compactness, which is reflected by having fewer parts (#Parts) and leads to ease of manipulation/editing. (3) Rendering quality, measured by the standard image quality metrics PSNR and SSIM.
@@ 5.1 Evaluation on ShapeNet
In particular, DPA-Net is better by 88%/17%/68% in CD/NC/ECD and uses 45% fewer 3D parts. Compared with the modified pixelNeRF, we obtain similar errors (differs by 0.3 in CD, 0.07 in NC, 0.8 in ECD, 0.55 in PSNR and 0.003 in SSIM) across all metrics. However, our method is more flexible in that it also produces convex part decomposition in an unsupervised manner, which is more useful in many applications, compared to the non-structured mesh produced by pixelNeRF.
@@
Method Input CD \!$\downarrow$ NC \!$\uparrow$ ECD \!$\downarrow$ #Parts \!$\downarrow$ PSNR \!$\uparrow$ SSIM \!$\uparrow$
EMS 3D 6.93 0.65 12.56 14.92 - -
pixelNeRF 2D 0.50 0.85 3.21 - 27.85 0.963
Ours 0.79 0.78 4.01 8.15 27.30 0.960
Ablation studies: We validate the effectiveness of four important strategies on improving DPA-Net: test-time adaptation (TTA), adaptive pixel sampling (AS), overlapping loss (OL) and primitive dropout (PD). Table 3 shows the main quantitative results. Removing TTA (w/o TTA) and adaptive sampling (w/o AS) significantly degrades the abstraction quality as the CD increases by 54% and 41% respectively, validating the contribution of our TTA strategy and silhouette-aware pixel sampling method. Removing the overlapping loss (w/o OL) results in higher number of primitives and convex shapes, showing that this loss helps in removing redundant, heavily overlapped convexes. The primitive dropout (PD) strategy further reduces the number of primitives by 29% without sacrificing the abstraction accuracy.
Fig. 4 qualitatively demonstrates the contributions of the four designs. TTA helps recover missing structures from the pretraining results, while adaptive pixel sampling aids in capturing thin structures and achieving accurate shape silhouettes. Enforcing a overlap loss removes redundant 3D parts. Primitive dropouts effectively reduce the number of primitives for assembling the final shape.
Figure 4: Ablation studies on the three key strategies to boost DPA-Net. The texts, #P’s, in the last column show the number of qudaric primitives used. The cyan ovals highlight challenging reconstruction areas that our method with TTA or adaptive sampling correctly recovers. The red ovals highlight redundant overlapping convex parts.
Table 3: Ablation studies on the ShapeNet chair benchmark. #P and #C denote the number of quadric primitives and convex shapes used during shape assembly. The colors cyan and orange represent the best and the second best results.
@@
More importantly, while pixelNeRF attained higher accuracy when trained solely on chairs, our DPA-Net surpasses pixelNeRF across all metrics (except for NC and PSNR) in this diverse multi-category experiment — for instance, improving on the CD metric by nearly 60%. This highlights the enhanced capacity of our structured primitive-based representation in capturing expansive shape variations compared to pixelNeRF’s unstructured approach.
@@
Method Input CD \!$\downarrow$ NC \!$\uparrow$ ECD \!$\downarrow$ #Parts \!$\downarrow$ PSNR \!$\uparrow$ SSIM \!$\uparrow$
EMS 3D 5.15 0.75 43.45 8.28 - -
pixelNeRF 2D 3.67 0.73 11.01 - 27.17 0.923
Ours 2D 1.47 0.70 4.91 5.57 25.96 0.932
@@ 5.2 Evaluation on DTU
As shown in Table 5, DPA-Net obtains higher-quality 3D abstractions with 18% lower average CD while using 3.5 more parts compared to DBW. DiViNet [46] and pixelNeRF [53] outperform DAP-Net as they output un-structured meshes. Fig. 6 provides visual comparisons showing that our method can recover fine structural details.
@@
Method Input Structure S24 S40 S55 S63 S83 S105 Mean CD Mean #Parts
EMS 3D 4.04 5.85 4.13 4.99 3.54 5.99 4.75 4.50
DBW 2D 3.85 1.34 3.83 4.57 5.05 6.07 4.12 3.83
Ours 2D 4.74 3.27 2.88 2.99 3.26 3.30 3.41 7.33
pixelNeRF 2D $\times$ 3.54 2.07 2.27 4.57 3.52 2.12 3.01 -
DiViNet 2D $\times$ 3.34 1.47 0.75 2.74 1.94 1.17 1.90 -
@@ 6 Conclusion
Acknowledgement
We thank all the anonymous reviewers and area chairs for their constructive comments. We also thank Kai Wang and Jianbo Ye both from Amazon for proofreading.